OpenStudy (anonymous):

2x^2+6x+4=0 factor it

OpenStudy (aum):

AC Method: a*c = 2*4 = 8 2 + 4 = 6 Therefore, 2x^2+6x+4=0 can be written as: 2x^2 + 2x + 4x + 4 = 0 2x(x + 1) + 4(x + 1) = 0 (2x + 4)(x + 1) = 0

OpenStudy (perl):

or more easily you can factor out 2 first

OpenStudy (perl):

2 (x^2 + 3x + 2 ) = 0

OpenStudy (aum):

true. (2x + 4)(x + 1) = 0 2(x+2)(x+1) = 0 (x+2)(x+1) = 0

OpenStudy (perl):

then you have 2(x+1)(x+2)

OpenStudy (anonymous):

ok got it

OpenStudy (aum):

If you are asked to factor it, leave it at 2(x+1)(x+2) = 0 If you are asked to solve for x, then (x+1)(x+2) = 0 and x = -1, -2

OpenStudy (anonymous):

ok